OpenStudy (anonymous):

Happy saturday everyone....need your brilliant help...:)))) Solve by substitution: x/2-y/3=5/6 x/5-y/4=-13/10

OpenStudy (amistre64):

id substitute a 6 and a 20

OpenStudy (anonymous):

x/2-y/3=5/6

OpenStudy (anonymous):

yup tried all of thos but my y and xs are ridiculous repetitive #

OpenStudy (experimentx):

x = (5/6+y/3)*2 ------------1 (5/6+y/3)*2/5 - y/4 = -13/10 ----2 on solving 2 y=14, from 1 x=11

OpenStudy (anonymous):

yowza thanks e...thanks for your help